OpenStudy (anonymous):

Composition of Functions: f(x) {(3,2),(5,2),(7,4),(9,3),(11,5)} g(x) {(1,3),(2,5),(3,7),(4,9),(5,11)}

OpenStudy (anonymous):

What is f(g(3)) ?

OpenStudy (anonymous):

@satellite73 @Hero

hero (hero):

Which do you believe is the correct answer?

OpenStudy (anonymous):

no i have to compose it

hero (hero):

I know what you have to do bro. I asked you a simple question to see, if perhaps, you have already attempted to solve it.

hero (hero):

With these ordered pairs, the first number is the input, the second number is the output.

OpenStudy (anonymous):

No actuallly i don't remember the method, nor can i find it in my textbook : /

hero (hero):

so what is the output of g(3)?

OpenStudy (anonymous):

is it 7?

hero (hero):

Yes, that is correct. So, in other words g(3) = 7

hero (hero):

So now you replace g(3) with 7: f(g(3)) = f(7) So what is the output of f(7) if x = 7 is the input?

OpenStudy (anonymous):

4 ?

hero (hero):

Yup, it was that easy.

OpenStudy (anonymous):

ohh, so now do i have to continue this for each one

hero (hero):

Good luck with it bro.

OpenStudy (anonymous):

@Hero (f-g)(x) how would i solve that ?

hero (hero):

(f - g)(x) = f(x) - g(x)